ARGOS Commissioning, June 28 to July 6, 2015

Players: ARGOS team

Report from Sebastian

  • Most work was closed dome using the calibration source. Quite some extense work has been conducted within this run.

  • In advance our mechanics team has sucesfully exchanged the SX pointing drives of the LM1 motors and gears in order to increase pointing accuracy of the laser guide stars. As well a heater system has been inserted behind LM2 on DX. Those measures now should be one of the last hardware actions we had to take in order to get the launch system more accurate and counteract the bending hextek mirrors. A test on sky and a setup of according control electronics is pending the next run. First many thanks to Matthias, Mathias and Franz for the good work!.

  • One of the major goals in this run was ranked around understanding the tip tilt control and APD vs, pyramid sensing schemes. With placing a filter and dimming the calibration source we could show that even with very low flux -estimated equal to 18..19 mag- the tilt and high order loop can be successfully closed in ~1.2'' ground layer turbulence placed on the secondary mirror. This result on the calibration source makes us very poitive that we will be able to achieve our goals as well on sky in next runs. The PSF size could be checked first this time on the LUCI N30, during day, showing nicely the expected performance.

  • Major work has been conducted on the control of the system: improving the sequence of aquisition mainly with the calibration source, pending more testing on sky. This is aimed at reaching more automatic operation for system calibration and operation.

  • In preparation for our future LAT- the 'Laser Alignment Telescope' that will aid a fully automatic LGS aquisition, Alex our new student at MPIA could take some LGS images on sky that will allow him to develop the required routines.

  • The one night we had with launched lasers, high cirrus made the closed loop testing difficult, but left us with quite some learning on field stop centering requirements and similar in high background wavefront sensing, but as well with moments of the highest flux on the wavefront sensors we have ever seen, when gating reaches into cirrus.

  • The next planned steps for ARGOS at LBT are coming late August with an installation run, bringing SX and upgraded DX sensors to the telescope, and two planned commissioning nighttime runs until christmas.

  • Many thanks to all the team, Gilles, Marco, Lorenzo, Martin, Jose, Wolfgang, Julian, Alex and of course Gustavo and Michael and the excellent LBTO support, spending days and nights with great enthusiasm here at LBT to proceed with ARGOS.

TBAD operation in parallel to Spotters

  • We were on-sky only one night, with more traffic than usual during the first half of the night, probably due to the 4th of July holiday.

  • Spotters/TBAD closures:
    • Night 1 (July 2 UTC): 7 closures called by spotters. 7 detected by TBAD (plus 3 before lasing).

Time (UTC) Spotter TBAD ID AZ EL Notes
03:09 - X A8233E 120.0 90.0 Not lasing. Spotters not out yet.
04:17 - X ACB132 107.5 54.0 Not lasing. Spotters not out yet.
04:26 - X A74F33 101.3 79.9 Not lasing. Spotters not out yet.
04:36 X X A7506A 47.5 50.9  
04:50 - X A82AAC 46.9 53.0 No activity reported on log.
05:18 X X AA7426 44.8 57.2  
05:23 X X AB6B02 44.3 57.9  
05:26 X X ACA9C4 44.0 58.3  
05:35 X X AD2833 42.9 59.7  
06:12 X - none 36.6 64.7 No TBAD closure.
09:02 - X AA8A07 -31.4 67.3 No closure reported on log. Only warning.
11:02 X - none 231.4 78.0 Probably spotter misidentification. No aircraft on map.

-- %USERSIG{GustavoRahmer - 2015-07-15}%

Comments

 
Topic revision: r1 - 15 Jul 2015, GustavoRahmer
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback